There is some evidence to suggest that businesses are moving out of states where unions are prevalent. In California, 18.4 % of all workers belong to a union, and in Arkansas, 3.7 % . Suppose 20 workers from California and 20 workers from Arkansas are selected at random. (Round your answers to four decimal places.) (a) Find the probability that at most three California workers belong to a union. (b) Find the probability that none of the Arkansas workers belongs to a union Find the probability that at least one worker from each state belongs to a union. (c)find probability that is at least one worker from each state belongs to a union
Answer #1

a) probability that at most three California workers belong to a union =P(X<=3)

= =0.4838

b)

probability that none of the Arkansas workers belongs to a union =(1-0.037)20 =0.4705

c)

probability that is at least one worker from each state belongs to a union =(1-(1-0.184)20)*(1-(1-0.037)20)=0.5205